OpenStudy (anonymous):

5.(15points) Letf(x)=e^−x^2. (a) What horizontal or vertical asymptotes does f(x) have? (b) Find the intervals of increase or decrease. (c) Find the local maximum and minimum values. (d) Where is f(x) concave downward? (e) Where are the inflection points? i have that th ederivatice is 2x e^(-x^2)

OpenStudy (anonymous):

it does not have vertical asymptotes since e^(x^2) is never zero

OpenStudy (anonymous):

u can find the horizontal asimptotes finding the limit when x goes to INF and u will see that the horizontal asymtota is y=0

OpenStudy (anonymous):

yeah i got that. I'm having a hard time finding the critical numbers other than 0

OpenStudy (anonymous):

it just have 0 like its critical number

OpenStudy (anonymous):

so for every negative x the derivative is <o

OpenStudy (anonymous):

the derivative is always negative so it is decreasing always except in x=0

OpenStudy (anonymous):

decreas = -infinity, 0 increase 0, infinity

OpenStudy (anonymous):

u forgot the (-2x) in the derivative

OpenStudy (anonymous):

ohright (-infinity, 0) and (0, infinity)

OpenStudy (anonymous):

wait there is a probkem i made a mistake

OpenStudy (anonymous):

the derivative is positive in a interval

OpenStudy (anonymous):

so was my first answer right?

OpenStudy (anonymous):

i know teh steps to solve this but what is throwing me off is the e^-x^2 because i don't know how this affects the fuction.

zepdrix (zepdrix):

\[\huge 0=-2xe^{-x^2}\] \[\large 0=-2x \qquad\rightarrow\qquad x=0\]\[\large 0=e^{-x^2} \qquad\rightarrow\qquad e^{-x^2}\ne0\] Yah you've got the right idea melo, x=0 is our only critical point. The exponential can never be zero. It just makes the function ...curvy.. and.. stuff :p

OpenStudy (anonymous):

ok what about my increasing and decreasing intervals

zepdrix (zepdrix):

So our critical point is at x=0. So the only place the function `can` change from increasing to decreasing or the other way around, is at x=0. So let's plug in a test point to see what's happening on the left side of x=0, and then on the right side as well. f'(-1)=?

OpenStudy (anonymous):

-2 e^-1

zepdrix (zepdrix):

Woops you missed an x there! D: \[\huge f'(-1)=-2(-1)e^{-1}\]

zepdrix (zepdrix):

We don't actually care about the `value` this is producing. We only care about it's `sign`. That being said, we can actually ignore the exponential part, since it will always be positive. So just plug in for the other part, and figure out the sign. \[\large f'(-1)=-2(-1)\]

OpenStudy (anonymous):

oh i missed the first negative sign actually -_- ok yes

zepdrix (zepdrix):

oh i see :o

OpenStudy (anonymous):

so the exponential sign is always positive.

OpenStudy (anonymous):

x<o is increasing and x>o is decreasing

zepdrix (zepdrix):

Yah that sounds right! Understand how to find inflection points?

OpenStudy (anonymous):

which mean at o it is a local max

zepdrix (zepdrix):

why are you using o? lol .. you mean 0? :p

OpenStudy (anonymous):

yes thats what i mean sorry x=0 is local max the second derivative is -2e^(-x^2)+xe^(-x^2)-2x right

zepdrix (zepdrix):

\[\large f''(x)=-2e^{-x^2}-2xe^{-x^2}(-2x)\]

zepdrix (zepdrix):

You apply the product rule. It looks like you did the first term correctly, simply differentiating the x giving you 1 in it's place. I'm not sure what you did beyond that though :o The next term should just give you another -2x, similar to the way you got your first derivative.

zepdrix (zepdrix):

\[\large f'(x)=-2xe^{-x^2}\] \[\large f''(x)=\color{royalblue}{\left(-2x\right)'}e^{-x^2}+(-2x)\color{royalblue}{\left(e^{-x^2}\right)'}\] \[\large f''(x)=\color{orangered}{\left(-2\right)}e^{-x^2}+(-2x)\color{orangered}{\left[e^{-x^2}(-2x)\right]}\] Understand how that works? :o

OpenStudy (anonymous):

yes sometimes i confuse myself when there is something in front of the x.I see my mistake. so to find the concavity with the second derivative i just need to see where it would be negative and positive. the inflection points im not sure.

zepdrix (zepdrix):

\[\large f(x)=0 \qquad\rightarrow\qquad\;\; \text{roots}\]\[\large f'(x)=0 \qquad\rightarrow\qquad \text{critical points}\]\[\large f''(x)=0 \qquad\rightarrow\qquad \text{inflection points}\]

zepdrix (zepdrix):

Critical points tell us where the function can change from increasing to decreasing. Inflection points can tell us where a function changes from CCU to CCD.

zepdrix (zepdrix):

Set the second derivative equal to zero. Find those inflection points! :) \[\large 0=-2e^{-x^2}+4x^2e^{-x^2}\]

OpenStudy (anonymous):

i dont know how to with the e's there

zepdrix (zepdrix):

See how the e's are the same in both terms? Factor it out, then we'll do similar step we did with the first derivative ~ Setting each factor equal to zero. \[\large 0=(-2+4x^2)e^{-x^2}\]

zepdrix (zepdrix):

We already determined earlier that \(\large e^{-x^2}\ne0\) right? So we only have to worry about the other factor,\[\large 0=-2+4x^2\]

OpenStudy (anonymous):

x= \[\sqrt{2}/2\]

zepdrix (zepdrix):

Careful, on the left side we had \(\large \sqrt{x^2}\) which will produce both a negative and positive solution.

zepdrix (zepdrix):

Or right side, depending on where you moved the stuff heh.

OpenStudy (anonymous):

so both the negative nd the positive are the inflection poins

zepdrix (zepdrix):

Yes, it appears we have 2 inflection points :)

OpenStudy (anonymous):

so lets say since at 0 the second derivative is negative then it is CCD

OpenStudy (anonymous):

in teh interval between the inflection points?

zepdrix (zepdrix):

yes, very good.

OpenStudy (anonymous):

thank you so much for the help

zepdrix (zepdrix):

np \c:/